At equal concentrations, would a nonelectrolyte (e.g. glucose) or electrolyte (e.g. NaCl) containing solution have lower freezing point? Why?
Answer #1

A solution containing an electrolyte will have a lower freezing point.

The depression in freezing point = i \times K_f \times m

where, i = number of ions in the solution, Kf = molal depression constant, m = molal concentration.

For any 2 solutions containg the same solvent and having same concentrations.

m is same and Kf is same.

The value of i for non-electrolyte = 1, the value of i for electrolyte >1.

example : glucose does not dissociate into ions (i = 1)

NaCl(aq) \rightarrow Na^+ (aq) + Cl^-(aq)   (i = 2)

because 2 ions = Na^+, Cl^- in solution

Thus,

For NON-ELECTROLYTE Solution, \Delta T = K_f \times m

For ELECTROLYTE Solution, \Delta T = i \times K_f \times m (i > 1)

Hence, depression in freezing point will be more for a solution containing an electrolyte.